# OpenLLM-Ro - Technical Report on Open-source Romanian LLMs trained starting from Llama 2

### v3

@@ 1 Introduction
We present and release111www.huggingface.co/OpenLLM-Ro/RoLlama2-7b-Base-v1 the first foundational Romanian LLM (i.e., RoLlama) based on the open-source Llama 2 model Touvron et al. (2023). Our collection of models is grouped under the OpenLLM-Ro initiative.
We collect and translate a large set of conversation and instructions, resulting in RoLlama2-7b-Instruct222www.huggingface.co/OpenLLM-Ro/RoLlama2-7b-Instruct-v1 and RoLlama2-7b-Chat333www.huggingface.co/OpenLLM-Ro/RoLlama2-7b-Chat-v1 models capable of following instructions and entertaining multi-turn conversations in Romanian.
@@ 4.2 Instruct & Chat Models
We train RoLlama2-7b-Instruct using translated instructions while we further include the translated conversations from UltraChat and OpenAssistant to the finetuning dataset for RoLlama2-7b-Chat. Training is done with a fixed learning rate of 5e-5 over one epoch.
@@ 5 Evaluation
Evaluating a general LLM is not a straightforward task considering the vast number of benchmarks and evaluation protocols proposed to judge the quality of a model. Popular benchmarks include Open LLM Leaderboard242424www.huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard, an online leaderboard that encompasses six reasoning and general knowledge tasks: ARC Clark et al. (2018), HellaSwag Zellers et al. (2019), MMLU Hendrycks et al. (2020), TruthfulQA Lin et al. (2021), Winogrande Sakaguchi et al. (2021) and GSM8k Cobbe et al. (2021). MT-Bench Zheng et al. (2024) proposes a multi-turn question set and shows that strong LLMs (e.g., GPT-4) can act as judges and can approximate human preferences in a scalable and somewhat explainable way. AGIEval Zhong et al. (2023) is a benchmark designed to evaluate foundational models in the context of human-centric standardized exams. Evaluation is done in multiple settings, from zero and few-shot to Chain of Thought (COT) prompting. While benchmarks in the Open LLM Leaderboard select the most probable answer (out of four candidates) for each question, the focus on MT-Bench is more on the quality of the generated text than the intrinsic knowledge of the world of each model. Some experiments252525www.twitter.com/maximelabonne/status/1746317053143773628?s=19 try and compute a correlation factor between human preferences and results of previously mentioned benchmarks.
Evaluating LLMs in any language besides English only adds another layer of complexity. For Romanian versions of ARC, HellaSwag, MMLU, and TruthfulQA, we resort to ChatGPT translations Dac Lai et al. (2023), while we translate262626same system used for translating instructions Winogrande and GSM8k. The generally used few-shot values for ARC, HellaSwag, and MMLU (i.e., 25 shots for ARC, 10 shots for HellaSwag, and 5 shots for MMLU) lead to the prompt having a length equal to or even surpassing 2048 tokens. As some of our models have been trained only on sequences of length 512, 1024, or 2048, we opt to use multiple few-shot settings. More specifically, for ARC, we test with 0, 1, 3, 5, 10, and 25 shots, MMLU with 0, 1, 3, and 5, HellaSwag with 0, 1, 3, 5, and 10 shots, and Winogrande is evaluated in 0, 1, 3 and 5 shot setting. For GSM8k, we employ 1, 3, and 5 shot evaluation. Results in Tables 1, 2 are averaged over a few shot settings for each benchmark. More detailed results are presented in the Appendix.
@@ 6 Results
We observe a very high performance of Mistral models (both foundational and chat). However, since the authors of Mistral (Jiang et al. (2023)) do not provide many details about their training data, we cannot know if the model has been trained on Romanian data as well as other languages. Furthermore, it seems that finetuning Mistral on Romanian Alpaca 272727www.huggingface.co/Andrei481/Mistral-7B-v0.1-Romanian/discussions/1 drastically decreases the performance on considered benchmarks.
Turning to an evaluation based more on generation rather than just picking the most probable continuation, the performance on MT-Bench is presented in Table 3. We note that even when using a Romanian prompt, most Llama2 and Mistral completions or answers are either entirely in English or have the first few words in Romanian, then reverting to English. Specifically asking the models to answer in Romanian in the prompt somewhat alleviates this issue, especially for Mistral models. As seen in the last column in Table 3, when specifically prompted to answer in Romanian, Llama2 manages to generate text in Romanian in only 25% of cases, while Mistral answers are almost entirely in Romanian (over 90% of cases vs 70% when not prompted to answer in Romanian). We automatically detect the language of the answer282828www.pypi.org/project/langdetect/ and overwrite with 0 the score of each answers that is not in Romanian.
As expected, our models can mostly generate text in Romanian without being specifically told do so. The chat model also generates 6 texts in English, which might be explained by the presence of some translation errors in the training conversations. With the exception of the Mistral model, the second-turn scores for all models are much lower than the first-turn scores. For RoLlama2-7b-Instruct, this is somewhat expected as the model has predominantly seen single-turn conversations during training. While still underperforming, the first turn score of RoLlama2-7b-Instruct is within 0.5 points of Mistral. Adding multi-turn conversations (i.e., in the form of UltraChat and Open Assistant) increases the score on second turns, but the overall score is still lower. The same phenomenon can also be observed in Table 1. As the conversations account for about 70% of the finetuning data (around 1.6M samples), their quality directly affects the performance of the final model. Methods for validating and curating the translations and selecting high-quality samples for building the best data blend should be researched further as they could lead to lower training times and better-performing models.
@@ 7 Conclusions
Furthermore, we present a general recipe which we expect to work on different, more powerful LLMs as well (e.g., bigger Llama2, Mistral, Mixtral292929www.mistral.ai/news/mixtral-of-experts/ or Llama3303030ai.meta.com/blog/meta-llama-3/ models). We firmly believe that this is just the first step towards building and adapting LLMs for Romanian for both research and industry applications.
@@ Acknowledgements
The universities contributed with researchers who worked pro-bono (POLITEHNICA Bucharest and the University of Bucharest) and computing power for training the models (POLITEHNICA Bucharest). Synchronization and collaboration took place within a project313131www.ilds.ro/llm-for-romanian carried out in the Institute of Logic and Data Science with funding from BRD Groupe Societe Generale.

### v4

@@ 1 Introduction
We present and release111www.huggingface.co/OpenLLM-Ro/RoLlama2-7b-Base the first foundational Romanian LLM (i.e., RoLlama) based on the open-source Llama 2 model Touvron et al. (2023). Our collection of models is grouped under the OpenLLM-Ro initiative.
We collect and translate a large set of conversation and instructions, resulting in RoLlama2-7b-Instruct222www.huggingface.co/OpenLLM-Ro/RoLlama2-7b-Instruct and RoLlama2-7b-Chat333www.huggingface.co/OpenLLM-Ro/RoLlama2-7b-Chat models capable of following instructions and entertaining multi-turn conversations in Romanian.
@@ 4.2 Instruct & Chat Models
We train RoLlama2-7b-Instruct using translated instructions while we further include the translated conversations from UltraChat and OpenAssistant to the finetuning dataset for RoLlama2-7b-Chat. Training is done with a fixed learning rate of 5e-5 over one epoch. We additionally train an instruct variant (with no pretraining on Romanian) of Mistral foundational model, leading to RoMistral-7b-Instruct242424www.huggingface.co/OpenLLM-Ro/RoMistral-7b-Instruct. Training a RoMistral foundational model coupled with improved Instruct and Chat variants will follow shortly.
@@ 5 Evaluation
Evaluating a general LLM is not a straightforward task considering the vast number of benchmarks and evaluation protocols proposed to judge the quality of a model. Popular benchmarks include Open LLM Leaderboard252525www.huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard, an online leaderboard that encompasses six reasoning and general knowledge tasks: ARC Clark et al. (2018), HellaSwag Zellers et al. (2019), MMLU Hendrycks et al. (2020), TruthfulQA Lin et al. (2021), Winogrande Sakaguchi et al. (2021) and GSM8k Cobbe et al. (2021). MT-Bench Zheng et al. (2024) proposes a multi-turn question set and shows that strong LLMs (e.g., GPT-4) can act as judges and can approximate human preferences in a scalable and somewhat explainable way. AGIEval Zhong et al. (2023) is a benchmark designed to evaluate foundational models in the context of human-centric standardized exams. Evaluation is done in multiple settings, from zero and few-shot to Chain of Thought (COT) prompting. While benchmarks in the Open LLM Leaderboard select the most probable answer (out of four candidates) for each question, the focus on MT-Bench is more on the quality of the generated text than the intrinsic knowledge of the world of each model. Some experiments262626www.twitter.com/maximelabonne/status/1746317053143773628?s=19 try and compute a correlation factor between human preferences and results of previously mentioned benchmarks.
Evaluating LLMs in any language besides English only adds another layer of complexity. For Romanian versions of ARC, HellaSwag, MMLU, and TruthfulQA, we resort to ChatGPT translations Dac Lai et al. (2023), while we translate272727same system used for translating instructions Winogrande and GSM8k. The generally used few-shot values for ARC, HellaSwag, and MMLU (i.e., 25 shots for ARC, 10 shots for HellaSwag, and 5 shots for MMLU) lead to the prompt having a length equal to or even surpassing 2048 tokens. As some of our models have been trained only on sequences of length 512, 1024, or 2048, we opt to use multiple few-shot settings. More specifically, for ARC, we test with 0, 1, 3, 5, 10, and 25 shots, MMLU with 0, 1, 3, and 5, HellaSwag with 0, 1, 3, 5, and 10 shots, and Winogrande is evaluated in 0, 1, 3 and 5 shot setting. For GSM8k, we employ 1, 3, and 5 shot evaluation. Results in Tables 1, 2 are averaged over a few shot settings for each benchmark. More detailed results are presented in the Appendix.
@@ 6 Results
We observe a very high performance of Mistral models (both foundational and chat). However, since the authors of Mistral (Jiang et al. (2023)) do not provide many details about their training data, we cannot know if the model has been trained on Romanian data as well as other languages. Furthermore, it seems that finetuning Mistral on Romanian Alpaca 282828www.huggingface.co/Andrei481/Mistral-7B-v0.1-Romanian/discussions/1 drastically decreases the performance on considered benchmarks.
Turning to an evaluation based more on generation rather than just picking the most probable continuation, the performance on MT-Bench is presented in Table 3. We note that even when using a Romanian prompt, most Llama2 and Mistral completions or answers are either entirely in English or have the first few words in Romanian, then reverting to English. Specifically asking the models to answer in Romanian in the prompt somewhat alleviates this issue, especially for Mistral models. As seen in the last column in Table 3, when specifically prompted to answer in Romanian, Llama2 manages to generate text in Romanian in only 25% of cases, while Mistral answers are almost entirely in Romanian (over 90% of cases vs 70% when not prompted to answer in Romanian). We automatically detect the language of the answer292929www.pypi.org/project/langdetect/ and overwrite with 0 the score of each answers that is not in Romanian.
As expected, our models can mostly generate text in Romanian without being specifically told to do so. The chat model also generates 6 texts in English, which might be explained by the presence of some translation errors in the training conversations. With the exception of the Mistral model, the second-turn scores for all models are much lower than the first-turn scores. For RoLlama2-7b-Instruct, this is somewhat expected as the model has predominantly seen single-turn conversations during training. While still underperforming, the first turn score of RoLlama2-7b-Instruct is within 0.5 points of Mistral. Adding multi-turn conversations (i.e., in the form of UltraChat and Open Assistant) increases the score on second turns, but the overall score is still lower. The same phenomenon can also be observed in Table 1. As the conversations account for about 70% of the finetuning data (around 1.6M samples), their quality directly affects the performance of the final model. Methods for validating and curating the translations and selecting high-quality samples for building the best data blend should be researched further as they could lead to lower training times and better-performing models. RoMistral-7b-Instruct further improves the quality of the first turn answer compared with Mistral-Instruct, becoming the best performing model. On second turn answers the difference is small. The Chat variant of RoMistral will probably outperform Mistral-Instruct in the quality of both turns.
@@ 7 Conclusions
Furthermore, we present a general recipe which we expect to work on different, more powerful LLMs as well (e.g., bigger Llama2, Mistral, Mixtral303030www.mistral.ai/news/mixtral-of-experts/ or Llama3313131ai.meta.com/blog/meta-llama-3/ models). We firmly believe that this is just the first step towards building and adapting LLMs for Romanian for both research and industry applications.
@@ Acknowledgements
The universities contributed with researchers who worked pro-bono (POLITEHNICA Bucharest and the University of Bucharest) and computing power for training the models (POLITEHNICA Bucharest). Synchronization and collaboration took place within a project323232www.ilds.ro/llm-for-romanian carried out in the Institute of Logic and Data Science with funding from BRD Groupe Societe Generale.